OpenStudy (anonymous):

The length of a diagonal of a square is 5 inches. Find the perimeter of the square

OpenStudy (anonymous):

Do you have the area of the square?

OpenStudy (anonymous):

Thats all the info I have

OpenStudy (anonymous):

This URL can help you out for sure http://www.algebra.com/algebra/homework/word/geometry/Geometry_Word_Problems.faq.question.131017.html

OpenStudy (anonymous):

This is the correct answer: To find the perimeter you have to find the length of the sides. Since it is a square, all the sides are equal. The diagonal=5. That divides the square in two right triangles.  By the Pythagorean theorem I get that \[a^2+b^2=c^2\] so c=5, a=x , and b=x therefore \[5^=x^2+x^2\rightarrow5=2x^2\rightarrow2.5=x^2\rightarrow x=1.58\]so the perimeter would be 4x=6.325 inches Hope I helped, lemme know by giving me a medal!

OpenStudy (anonymous):

6.325 is incorrect. Pytjagoras sez that the square of the hypotenuse equals the sum of the squares of the sides. So "x" would be: sqrt((5^2)/2) The perimeter is then 4x

OpenStudy (anonymous):

12.25?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

What's 5^2?

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Oh i forgot to square it. What a huge typo lol. The answer is actually \[5^2=2x^2\rightarrow x=3.536\] Therefore the perimeter would be 4x=14.14 inches

OpenStudy (anonymous):

Thanks @qweqwe123123123123111

OpenStudy (anonymous):

NP, @4kec4 ! 3.356 is more better... :-)

OpenStudy (anonymous):

You hear that @adillie ? you should get 14.142 for your perimeter

OpenStudy (anonymous):

@qweqwe123123123123111 its getting late, maybe i should get some sleep seeing that i made that typo lol

OpenStudy (anonymous):

Heheh... This error pales next to some of the doozies I've made on here!

OpenStudy (anonymous):

sorry internet connection was lost, 25

OpenStudy (anonymous):

Thank you both sorry I couldnt reply!

OpenStudy (anonymous):

no problem anytime